# Both Stellar Mass and Gravitational Potential Shape the Gas-Phase Metallicity

### v1

@@ 4.1.4 The role of SFR
As previously mentioned, we do find that the SFR plays a role in determining metallicity, consistent with expectations from the FMR. This role does not vary significantly as a function of the radius within which the metallicity is estimated. If the dependence on SFR is driven by gas accretion, which both dilutes the metallicity and boosts the star formation, then this would imply that this process is ongoing on all galactic scales. This would be in line with independent, detailed studies of extraplanar gas kinematics in some nearby galaxies (e.g. Li et al., 2021, 2023).
It is, however, somewhat puzzling that the importance of the SFR is significantly lower than the $\sim 30\%$ originally estimated by the FMR (Mannucci et al., 2010; Curti et al., 2020). Using MaNGA metallicities within $\rm 1R_{e}$, the metallicity importance is only 7% (Fig.1). At $\rm 0.5R_{e}$ the importance of SFR increases slightly, but it is still only about 15% (Fig.2). Note that this is regardless of whether we use the total SFR in the galaxy (yellow curve in Fig.2), or the SFR estimated within a given radius (orange curve in Fig.2). Even in the full SDSS sample, the role of the SFR is still below 20% (Fig.3). A likely explanation is that in the original FMR works, where only $\mathrm{M_{*}}$ and SFR were considered, the SFR indirectly picked the metallicity dependence on other galaxy parameters that were not included in the original analysis, such as $\Phi$ and $\sigma$, which are instead considered in these RF studies. This is also supported by the fact that the importance of $\mathrm{M_{*}}$, at the level of $\sim$70%, is instead consistent with the original FMR studies.
@@ 4.2.1 The dependence on the stellar gravitational potential
Fig. 5 illustrates the results, including the signs, of the PCCs between metallicities and various galactic quantities in the MaNGA sample. As usual, the metallicity shows the strongest (positive) correlation with stellar mass. As discussed, $\mathrm{\Phi_{*}}$ cannot be included here, because it contains $\rm M_{*}$. However, here we are showing $\mathrm{M_{dyn,R_{e}}}$ and $\mathrm{\Phi_{dyn,R_{e}}}$, which we recall are calculated within $\rm 1R_{e}$. The metallicity shows a secondary dependence on these quantities; however, on the contrary to what was expected from the very positive dependence on $\mathrm{M_{*}}$ and $\mathrm{\Phi_{*}}$ discussed above (and shown in Fig.4), the dependence on $\mathrm{M_{dyn,R_{e}}}$ and $\mathrm{\Phi_{dyn,R_{e}}}$ are negative. We explore this effect more closely in Fig.6, where we show the 2D plot of the distribution of galaxies binned in $\mathrm{\Phi_{dyn,R_{e}}}$ and $\mathrm{\Phi_{*}}$, colour coded by metallicity within $\rm 1R_{e}$ and $2\rm R_{e}$. This diagram clearly illustrates the much stronger positive dependence on $\mathrm{\Phi_{*}}$ relative to $\mathrm{\Phi_{dyn},R_{e}}$ in both cases. When controlling for $\mathrm{\Phi_{*}}$, the dependence on $\mathrm{\Phi_{dyn,R_{e}}}$ for the metallicity within $\rm 1R_{e}$ is positive, while null for the metallicity within $\rm 2R_{e}$. This striking difference between the behaviour relative to $\mathrm{\Phi_{dyn,R_{e}}}$ and $\mathrm{\Phi_{*}}$ is puzzling, as they should both be tracing the gravitational potential. While in the representation of Fig.6 $\mathrm{\Phi_{*}}$ may have a strong contribution from the positive dependence on $\mathrm{M_{*}}$, that cannot be the entire explanation, given that we have seen in Fig.4 and in the previous section, that there is also a strong direct and casual dependence on $\mathrm{\Phi_{*}}$. Additionally, we have seen in Fig.5 that there is a negative metallicity correlation also with $\mathrm{M_{dyn,R_{e}}}$, regardless of $\mathrm{\Phi_{dyn,R_{e}}}$. The key to understanding the puzzle probably is in the fact that $\mathrm{\Phi_{dyn,R_{e}}}$ is determined within $\rm 1R_{e}$, while $\mathrm{\Phi_{*}}$ is determined by taking the whole mass of the galaxy, even well outside $\rm 2R_{e}$. Therefore, $\mathrm{\Phi_{*}}$ is probably a proxy of the global gravitational potential of the galaxy, while $\mathrm{\Phi_{dyn,R_{e}}}$ is only tracing the gravitational potential within the inner ($\rm 1R_{e}$) region. Therefore, the strongly positive metallicity correlation with $\mathrm{\Phi_{*}}$ may indeed reflect the global capability of the galaxy in retaining metals. The inverse dependence on $\mathrm{\Phi_{dyn,R_{e}}}$ is more difficult to understand, but has also been found by a much smaller sample in Baker & Maiolino (2023) (see Fig. 2 of that work). We speculate that a deeper inner gravitational potential may be the sink for a stronger inflow of metal-poor, or nearly pristine, gas towards the central region of the galaxy, which would dilute the metallicity, hence resulting in a negative correlation. This would be a local, more subtle version of a similar effect seen more prominently at high-z, in the form of inverted metallicity gradients (Cresci et al., 2010; Troncoso et al., 2014).
@@ 4.2.2 The dependence on SFR
We finally discuss the dependence on the SFR. In section 4.1.4, the RF confirmed a dependence on SFR, which is actually of third importance after accounting for $\mathrm{M_{*}}$ and $\Phi_{*}$. Here we explore the sign of the dependence by leveraging the PCC information. Fig.5-right shows a strong negative metallicity dependence (i.e. anti-correlation) with SFR, as expected from the FMR. However, this is true only for the metallicity estimated within $\rm 2R_{e}$ (right panel of Fig.5). When considering the metallicity within $\rm 1R_{e}$ (left panel of Fig.5), the sign of the correlation for the full sample becomes positive, i.e. opposite of what is expected from the FMR. We explore this phenomenon further in the 2D plot of Fig.7, which shows the distribution of MaNGA galaxies in the SFR versus $\rm M_{*}$ diagram, colour coded by metallicity within $\rm 1R_{e}$. The inverted dependence is driven by the very massive galaxies, which is where the mass-metallicity relation saturates, or becomes even slightly inverted, as already pointed out in the past by other authors (Yates et al., 2012; Peng & Maiolino, 2014a; Duarte Puertas et al., 2022); this is because the metallicity reaches the effective yield, beyond which additional star formation or mass growth does not significantly raise the metallicity (Tremonti et al., 2004; Peng & Maiolino, 2014a). The possible inversion of the FMR has also been attributed to gradual dilution following gas-rich mergers (Yates et al., 2012), and environmentally enriched inflows in satellites residing in dense regions (Peng & Maiolino, 2014a). If we restrict the sample to galaxies less massive than $\mathrm{\log{(M_{*}/M_{\odot})}<10.5}$, then the dependence on the SFR is back to the negative value expected from the FMR; this is illustrated by the magenta arrow angle in Fig.7 and by the magenta histogram in Fig.5. Notably, excluding high mass galaxies from our sample also results in significantly lower errors for the arrow angle.
@@ 5 Discussion
One of the most important results is that stellar mass $M_{*}$ and stellar gravitational potential $\Phi_{*}$ (determined through the spectro-photometric method) are the most important parameters in determining the gas metallicity in galaxies. However, their relative importance swaps as a function of the radius within which the metallicity is determined. In the central region ($1~R_{e}$) stellar mass is the most important parameter determining the metallicity, while gravitational potential plays a minimal role. Our interpretation is that in the central region the potential well is always deep enough to retain metals, regardless of the large-scale gravitational potential. As a consequence, the amount of metals simply reflects the integrated production of stars, i.e. the stellar mass. On large scales ($1.5R_{e}$), gravitational potential (as traced by $\Phi_{*}$) is the most important parameter driving the metallicity. In such outer regions, the gravitational potential likely plays a prominent role in the retention of metals.
The role of the SFR (the third most important parameter in determining the metallicity) does not have a strong dependence on the spatial scale, especially when considering galaxies with masses smaller than $10^{10.5}M_{\odot}$. If the metallicity’s inverse dependence on the SFR is interpreted as dilution from accreting pristine, or near-pristine gas, then our result may indicate that such accretion happens at the same level across the entire disc, as indeed observed in detailed studies of some individual nearby galaxies (Ho et al., 2015; Kreckel et al., 2019, 2025). It is also a very interesting finding that, at any scale, even if it is the third most important parameter, the SFR contributes only at the 7–15% level to determining the gas metallicity, which is much less than estimated by previous studies on the FMR (Mannucci et al., 2010; Curti et al., 2020). The likely origin of this difference is that most of the previous studies had only considered stellar mass and SFR, and not the several other parameters considered here (gravitational potential, velocity dispersion, size, etc.); as a consequence, in those previous studies SFR may have picked the importance in determining the metallicity from the parameters that were not included in their analysis, and which probably correlate with the SFR.
Finally, the finding that dynamical mass or total gravitational potential, as inferred from the dynamical mass, contribute very little to the metallicity of the galaxy, can be understood in terms of these quantities being derived only within $1R_{e}$. The fact that these dynamical tracers within $1R_{e}$ play little role in determining the metallicity further supports the scenario in which within the central region galaxies chemically evolve without loss of metals (or manage to reaccrete them locally), regardless of the depth of the gravitational potential in these regions.
@@ 6 Conclusions
  • We find that the primary drivers of the metallicity are stellar mass, $\mathrm{M_{*}}$, and the stellar gravitational potential, $\mathrm{\Phi_{*}=M_{*}/R_{e}}$ (which is likely a proxy of the total gravitational potential). However, we find that their relative role has a strong radial dependence: at small galactocentric distances ($\mathrm{\lesssim 1.25R_{e}}$), metallicity is primarily driven by stellar mass, whereas at larger radii ($\mathrm{\gtrsim 1.25R_{e}}$), $\mathrm{\Phi_{*}}$ becomes the dominant factor. We attribute this to the fact that central regions of galaxies experience rapid metal enrichment and efficient retention due to short chemical timescales and deeper gravitational wells, while in the outskirts of galaxies, chemical mixing is less efficient and gravitational potential is more influential in regulating outflows (hence metal retention) and, possibly, also inflows (hence metals dilution).
@@ Acknowledgements
We thank Mirko Curti for kindly providing the code used in our gas-phase metallicity measurements.
MK thanks the University of Cambridge Harding Distinguished Postgraduate Scholars Programme, UK Science and Technology Facilities Council (STFC) Center for Doctoral Training (CDT) in Data Intensive Science, and Girton College Cambridge for a PhD studentship.
RM acknowledges support from the Science and Technology Facilities Council (STFC), by the European Research Council (ERC) through Advanced Grant 695671 “QUENCH”, by the UK Research and Innovation (UKRI) Frontier Research grant RISEandFALL. RM also acknowledges support from a Royal Society Research Professorship grant.
WMB gratefully acknowledges support from DARK via the DARK fellowship. This work was supported by a research grant (VIL54489) from VILLUM FONDEN.
Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss4.org.
SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian (CfA), the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatório Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
@@
The SDSS DR7 data are publicly available. The MaNGA data used are also publicly available at https://www.sdss4.org/dr17/manga/manga-data/. The dynamical masses and velocity dispersions are available from Zhu et al. (2023). The MPA-JHU catalogue is publicly available at https://wwwmpa.mpa-garching.mpg.de/SDSS/DR7/. The morphological parameters of SDSS galaxies are available from Simard et al. (2011). The satellite classifications from Yang et al. (2005, 2007) are available at https://gax.sjtu.edu.cn/data/Group.html.

### v2

@@ 4.1.4 The role of SFR
As previously mentioned, we do find that the SFR plays a role in determining metallicity, consistent with expectations from the FMR. This role does not vary significantly as a function of the radius within which the metallicity is estimated. If the dependence on SFR is driven by gas accretion, which both dilutes the metallicity and boosts the star formation, then this would imply that this process is ongoing on all galactic scales. This would be in line with independent, detailed studies of extraplanar gas kinematics in some nearby galaxies (e.g. li_kinematic_2021; li_fountain-driven_2023).
It is, however, somewhat puzzling that the importance of the SFR is significantly lower than the $\sim 30\%$ originally estimated by the FMR (mannucci_fundamental_2010; curti_mass-metallicity_2020). Using MaNGA metallicities within $\rm 1R_{e}$, the metallicity importance is only 7% (Fig.1). At $\rm 0.5R_{e}$ the importance of SFR increases slightly, but it is still only about 15% (Fig.2). Note that this is regardless of whether we use the total SFR in the galaxy (yellow curve in Fig.2), or the SFR estimated within a given radius (orange curve in Fig.2). Even in the full SDSS sample, the role of the SFR is still below 20% (Fig.3). A likely explanation is that in the original FMR works, where only $\mathrm{M_{*}}$ and SFR were considered, the SFR indirectly picked the metallicity dependence on other galaxy parameters that were not included in the original analysis, such as $\Phi$ and $\sigma$, which are instead considered in these RF studies. This is also supported by the fact that the importance of $\mathrm{M_{*}}$, at the level of $\sim$70%, is instead consistent with the original FMR studies.
@@ 4.2.1 The dependence on the stellar gravitational potential
Fig. 5 illustrates the results, including the signs, of the PCCs between metallicities and various galactic quantities in the MaNGA sample. As usual, the metallicity shows the strongest (positive) correlation with stellar mass. As discussed, $\mathrm{\Phi_{*}}$ cannot be included here, because it contains $\rm M_{*}$. However, here we are showing $\mathrm{M_{dyn,R_{e}}}$ and $\mathrm{\Phi_{dyn,R_{e}}}$, which we recall are calculated within $\rm 1R_{e}$. The metallicity shows a secondary dependence on these quantities; however, on the contrary to what was expected from the very positive dependence on $\mathrm{M_{*}}$ and $\mathrm{\Phi_{*}}$ discussed above (and shown in Fig.4), the dependence on $\mathrm{M_{dyn,R_{e}}}$ and $\mathrm{\Phi_{dyn,R_{e}}}$ are negative. We explore this effect more closely in Fig.6, where we show the 2D plot of the distribution of galaxies binned in $\mathrm{\Phi_{dyn,R_{e}}}$ and $\mathrm{\Phi_{*}}$, colour coded by metallicity within $\rm 1R_{e}$ and $2\rm R_{e}$. This diagram clearly illustrates the much stronger positive dependence on $\mathrm{\Phi_{*}}$ relative to $\mathrm{\Phi_{dyn},R_{e}}$ in both cases. When controlling for $\mathrm{\Phi_{*}}$, the dependence on $\mathrm{\Phi_{dyn,R_{e}}}$ for the metallicity within $\rm 1R_{e}$ is positive, while null for the metallicity within $\rm 2R_{e}$. This striking difference between the behaviour relative to $\mathrm{\Phi_{dyn,R_{e}}}$ and $\mathrm{\Phi_{*}}$ is puzzling, as they should both be tracing the gravitational potential. While in the representation of Fig.6 $\mathrm{\Phi_{*}}$ may have a strong contribution from the positive dependence on $\mathrm{M_{*}}$, that cannot be the entire explanation, given that we have seen in Fig.4 and in the previous section, that there is also a strong direct and casual dependence on $\mathrm{\Phi_{*}}$. Additionally, we have seen in Fig.5 that there is a negative metallicity correlation also with $\mathrm{M_{dyn,R_{e}}}$, regardless of $\mathrm{\Phi_{dyn,R_{e}}}$. The key to understanding the puzzle probably is in the fact that $\mathrm{\Phi_{dyn,R_{e}}}$ is determined within $\rm 1R_{e}$, while $\mathrm{\Phi_{*}}$ is determined by taking the whole mass of the galaxy, even well outside $\rm 2R_{e}$. Therefore, $\mathrm{\Phi_{*}}$ is probably a proxy of the global gravitational potential of the galaxy, while $\mathrm{\Phi_{dyn,R_{e}}}$ is only tracing the gravitational potential within the inner ($\rm 1R_{e}$) region. Therefore, the strongly positive metallicity correlation with $\mathrm{\Phi_{*}}$ may indeed reflect the global capability of the galaxy in retaining metals. The inverse dependence on $\mathrm{\Phi_{dyn,R_{e}}}$ is more difficult to understand, but has also been found by a much smaller sample in baker_stellar_2023 (see Fig. 2 of that work). We speculate that a deeper inner gravitational potential may be the sink for a stronger inflow of metal-poor, or nearly pristine, gas towards the central region of the galaxy, which would dilute the metallicity, hence resulting in a negative correlation. This would be a local, more subtle version of a similar effect seen more prominently at high-z, in the form of inverted metallicity gradients (cresci_gas_2010; troncoso_metallicity_2014).
@@ 4.2.2 The dependence on SFR
We finally discuss the dependence on the SFR. In section 4.1.4, the RF confirmed a dependence on SFR, which is actually of third importance after accounting for $\mathrm{M_{*}}$ and $\Phi_{*}$. Here we explore the sign of the dependence by leveraging the PCC information. Fig.5-right shows a strong negative metallicity dependence (i.e. anti-correlation) with SFR, as expected from the FMR. However, this is true only for the metallicity estimated within $\rm 2R_{e}$ (right panel of Fig.5). When considering the metallicity within $\rm 1R_{e}$ (left panel of Fig.5), the sign of the correlation for the full sample becomes positive, i.e. opposite of what is expected from the FMR. We explore this phenomenon further in the 2D plot of Fig.7, which shows the distribution of MaNGA galaxies in the SFR versus $\rm M_{*}$ diagram, colour coded by metallicity within $\rm 1R_{e}$. The inverted dependence is driven by the very massive galaxies, which is where the mass-metallicity relation saturates, or becomes even slightly inverted, as already pointed out in the past by other authors (yates_relation_2012; peng_dependence_2014; duarte_puertas_mass-metallicity_2022); this is because the metallicity reaches the effective yield, beyond which additional star formation or mass growth does not significantly raise the metallicity (tremonti_origin_2004; peng_dependence_2014). The possible inversion of the FMR has also been attributed to gradual dilution following gas-rich mergers (yates_relation_2012), and environmentally enriched inflows in satellites residing in dense regions (peng_dependence_2014). If we restrict the sample to galaxies less massive than $\mathrm{\log{(M_{*}/M_{\odot})}<10.5}$, then the dependence on the SFR is back to the negative value expected from the FMR; this is illustrated by the magenta arrow angle in Fig.7 and by the magenta histogram in Fig.5. Notably, excluding high mass galaxies from our sample also results in significantly lower errors for the arrow angle.
@@ 5 Discussion
One of the most important results is that stellar mass $\rm M_{*}$ and stellar gravitational potential $\rm\Phi_{*}$ (determined through the spectro-photometric method) are the most important parameters in determining the gas metallicity in galaxies. However, their relative importance swaps as a function of the radius within which the metallicity is determined. In the central region ($\rm<1~R_{e}$) stellar mass is the most important parameter determining the metallicity, while gravitational potential plays a minimal role. Our interpretation is that in the central region the potential well is always deep enough to retain metals, regardless of the large-scale gravitational potential. As a consequence, the amount of metals simply reflects the integrated production of stars, i.e. the stellar mass. On large scales ($\rm>1.5R_{e}$), gravitational potential (as traced by $\rm\Phi_{*}$) is the most important parameter driving the metallicity. In such outer regions, the gravitational potential likely plays a prominent role in the retention of metals.
The role of the SFR (the third most important parameter in determining the metallicity) does not have a strong dependence on the spatial scale, especially when considering galaxies with masses smaller than $\rm 10^{10.5}M_{\odot}$. If the metallicity’s inverse dependence on the SFR is interpreted as dilution from accreting pristine, or near-pristine gas, then our result may indicate that such accretion happens at the same level across the entire disc, as indeed observed in detailed studies of some individual nearby galaxies (ho_metallicity_2015; kreckel_mapping_2019; kreckel_temperature-based_2025). It is also a very interesting finding that, at any scale, even if it is the third most important parameter, the SFR contributes only at the 7–15% level to determining the gas metallicity, which is much less than estimated by previous studies on the FMR (mannucci_fundamental_2010; curti_mass-metallicity_2020). The likely origin of this difference is that most of the previous studies had only considered stellar mass and SFR, and not the several other parameters considered here (gravitational potential, velocity dispersion, size, etc.); as a consequence, in those previous studies SFR may have picked the importance in determining the metallicity from the parameters that were not included in their analysis, and which probably correlate with the SFR.
Finally, the finding that dynamical mass or total gravitational potential, as inferred from the dynamical mass, contribute very little to the metallicity of the galaxy, can be understood in terms of these quantities being derived only within $\rm 1R_{e}$. The fact that these dynamical tracers within $\rm 1R_{e}$ play little role in determining the metallicity further supports the scenario in which within the central region galaxies chemically evolve without loss of metals (or manage to reaccrete them locally), regardless of the depth of the gravitational potential in these regions.
@@ 6 Conclusions
  • We find that the primary drivers of the metallicity are stellar mass, $\mathrm{M_{*}}$, and the stellar gravitational potential, $\mathrm{\Phi_{*}=M_{*}/R_{e}}$ (which is likely a proxy of the total gravitational potential). However, we find that their relative role has a strong radial dependence: at small galactocentric distances ($\mathrm{\lesssim 0.7R_{e}}$), metallicity is primarily driven by stellar mass, whereas at larger radii ($\mathrm{\gtrsim 0.9R_{e}}$), $\mathrm{\Phi_{*}}$ becomes the dominant factor. We attribute this to the fact that central regions of galaxies experience rapid metal enrichment and efficient retention due to short chemical timescales and deeper gravitational wells, while in the outskirts of galaxies, chemical mixing is less efficient and gravitational potential is more influential in regulating outflows (hence metal retention) and, possibly, also inflows (hence metals dilution).
@@ Acknowledgements
We are grateful to the referee, Jianhui Lian, for helpful comments that have improved the manuscript. We thank Mirko Curti for kindly providing the code used in our gas-phase metallicity measurements. We are very grateful for feedback and comments from Francesco Belfiore and Asa Bluck. MK thanks the University of Cambridge Harding Distinguished Postgraduate Scholars Programme, UK Science and Technology Facilities Council (STFC) Center for Doctoral Training (CDT) in Data Intensive Science, and Girton College Cambridge for a PhD studentship. RM acknowledges support from the Science and Technology Facilities Council (STFC), by the European Research Council (ERC) through Advanced Grant 695671 “QUENCH”, by the UK Research and Innovation (UKRI) Frontier Research grant RISEandFALL. RM also acknowledges support from a Royal Society Research Professorship grant. WMB gratefully acknowledges support from DARK via the DARK fellowship. This work was supported by a research grant (VIL54489) from VILLUM FONDEN. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian (CfA), the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatório Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
@@
The SDSS DR7 data are publicly available. The MaNGA data used are also publicly available at https://www.sdss4.org/dr17/manga/manga-data/. The dynamical masses and velocity dispersions are available from zhu_manga_2023. The MPA-JHU catalogue is publicly available at https://wwwmpa.mpa-garching.mpg.de/SDSS/DR7/. The morphological parameters of SDSS galaxies are available from simard_catalog_2011. The satellite classifications from yang_halo-based_2005; yang_galaxy_2007 are available at https://gax.sjtu.edu.cn/data/Group.html.